OpenStudy (anonymous):

What is the equation for the axis of symmetry in the function? Write the equation as x=__, filling in the blank. y=-2x^2+5x-4

OpenStudy (campbell_st):

the axis of symmetry is found using x = -b/2a in your question you have b = 5 and a = -2 so x = -5/(2*-2) x = 5/4

OpenStudy (anonymous):

thnx